OpenStudy (anonymous):

walk me through this problem: 5x+30/x+4 x x+4/x^2+12x+36

OpenStudy (luigi0210):

Multiplying fractions?

OpenStudy (anonymous):

I would suppose

OpenStudy (luigi0210):

Let's start by writing it out so we can see more clearly: \[\LARGE \frac{5x+30}{x+4}*\frac{x+4}{x^2+12x+36}\]

OpenStudy (luigi0210):

First step, factor what you can.

OpenStudy (anonymous):

factor x+4??

OpenStudy (luigi0210):

Well it's already factored, so move on to what you can factor. Try the 5x+30 and x^2+12x+36

OpenStudy (anonymous):

you can factor the 5 through 5x+30?

OpenStudy (luigi0210):

Yup, so what do you get?

OpenStudy (anonymous):

x+6

OpenStudy (luigi0210):

\[\LARGE \frac{5(x+6)}{x+4}*\frac{x+4}{x^2+12x+36}\] Yup, now move on to the next

OpenStudy (anonymous):

factor the 12 through x^2+12x+36 and get x^2+x+3?

OpenStudy (luigi0210):

If there was a 12 infront of the x^2 it would work, but not in this case.

OpenStudy (anonymous):

then idk :l

OpenStudy (luigi0210):

Well, what adds up to 12 and multiplies to 36?

OpenStudy (luigi0210):

Factor 36.

OpenStudy (anonymous):

6

OpenStudy (luigi0210):

Right! So it would just be (x+6)(x+6) or the easier way to write it is (x+6)^2. \[\LARGE \frac{5(x+6)}{x+4}*\frac{x+4}{(x+6)^2}\] Now that we've factored, we can cross cancel anything that's the same. So notice anything that's the same?

OpenStudy (anonymous):

(x+6) and x+4

OpenStudy (luigi0210):

Yup, they cancel and we are left with: \[\LARGE \frac{5}{1}*\frac{1}{(x+6)}\] Now just multiply that :)

OpenStudy (anonymous):

which would be 5x+30?

OpenStudy (luigi0210):

If it was \[\LARGE \frac{1}{5}*\frac{1}{x+6}\] then yes, but not in this case. We multiply across, so: \[\LARGE \frac{5*1}{1*(x+6)}=\frac{5}{x+6}\]

OpenStudy (luigi0210):

Two completely different answers, so be careful.

OpenStudy (anonymous):

ahh thank you

OpenStudy (luigi0210):

Anytime :)